# Finch members

### v2

@@ 4 Related Work
Recent years have seen significant progress in benchmarks for financial reasoning [chen2021finqa, islam2023financebench, finance_agent_benchmark, finben2024, findabench2025, finagentbench2025, finauditing2025, li2025investorbench, visfineval2025, gdpval2025, finsearchcomp2025, xfinbench2025], spreadsheet reasoning [li2023sheetcopilot, chen2024sheetagent, ma2024spreadsheetbench, sodbench2025, li2023auto, dong2019tablesense, sheetmind2025, spreadsheetllm2024, wu2025realhitbench, li2025mimotable, dong2025machinelearninglm, dong2025reasoning, li2024tapilot, zhao2024nl2formula, odysseybench2025, officebench2024], and multimodal document [mathew2021docvqa], table [li2020tablebank, zheng2024multimodal], chart [wang2024charxiv], table-chart [dong2024ttc], and spreadsheet [xia2024vision, dong2019tablesense] reasoning, driving advances in LLM-based agents for enterprise tasks. However, Finch proposes a new benchmark for the messy artifacts and long-horizon workflows in wild F&A enterprise settings.
@@ 5 Conclusion
We introduced Finch, a new benchmark for real-world F&A enterprise workflows. Finch combines workflows induced from enterprise email threads, version histories of spreadsheets, and high-quality financial artifacts with rigorous expert annotation and a calibrated LLM-as-judge framework, enabling systematic evaluation of agents on diverse workflows that operate over large, messy, and multimodal enterprise artifacts and require long-horizon, spreadsheet-centric reasoning. Our experiments show that even the strongest frontier systems pass fewer than 40% of workflows after spending 48 hours on the benchmark, revealing a substantial gap between current AI capabilities and the demands of real enterprise practice. We hope Finch will serve as a foundation for developing agents to tackle real, messy and long-horizon professional work.
@@ Appendix C Detailed Analysis
Our API-based runs are single-call: they leverage the models’ underlying reasoning capabilities but lack two crucial affordances that web agents exploit. (i) interleaved code execution with feedback, and (ii) explicit reflection based on intermediate tool outputs. As a result, the API agents must generate the entire plan, code, and outputs within a single LLM call. When their initial structural assumptions about a spreadsheet are slightly off, they have no mechanism to detect or correct the mistake, leading to a significantly higher error rate, particularly in categories related to schema understanding and table manipulation. It’s desirable for future work to explore agentic methods with multiple round of API calls.

### v3

@@ 4 Related Work
Recent years have seen significant progress in benchmarks for financial reasoning [chen2021finqa, islam2023financebench, finance_agent_benchmark, finben2024, findabench2025, finagentbench2025, finauditing2025, li2025investorbench, visfineval2025, gdpval2025, finsearchcomp2025, xfinbench2025, xie2024finben], spreadsheet reasoning [li2023sheetcopilot, chen2024sheetagent, ma2024spreadsheetbench, sodbench2025, li2023auto, dong2019tablesense, sheetmind2025, spreadsheetllm2024, wu2025realhitbench, li2025mimotable, dong2025machinelearninglm, dong2025reasoning, li2024tapilot, zhao2024nl2formula, odysseybench2025, officebench2024], and multimodal document [mathew2021docvqa], table [li2020tablebank, zheng2024multimodal], chart [wang2024charxiv], table-chart [dong2024ttc], and spreadsheet [xia2024vision, dong2019tablesense] reasoning, driving advances in LLM-based agents for enterprise tasks. However, Finch proposes a new benchmark for the messy artifacts and long-horizon workflows in wild F&A enterprise settings.
@@ 5 Conclusion
We introduced Finch, a new benchmark for real-world F&A enterprise workflows. Finch combines workflows induced from enterprise email threads, version histories of spreadsheets, and high-quality financial artifacts with rigorous expert annotation and a calibrated LLM-as-judge framework, enabling systematic evaluation of agents on diverse workflows that operate over large, messy, and multimodal enterprise artifacts and require long-horizon, spreadsheet-centric reasoning. Our experiments show that even the strongest frontier systems pass fewer than 40% of workflows after spending 16.8 minutes per workflow, revealing a substantial gap between current AI capabilities and the demands of real enterprise practice. We hope Finch will serve as a foundation for developing agents to tackle real, messy and long-horizon professional work.
@@ Appendix C Detailed Analysis
Our API-based runs are single-call: they leverage the models’ underlying reasoning capabilities but lack two crucial affordances that web agents exploit. (i) interleaved code execution with feedback, and (ii) explicit reflection based on intermediate tool outputs. As a result, the API agents must generate the entire plan, code, and outputs within a single LLM call. When their initial structural assumptions about a spreadsheet are slightly off, they have no mechanism to detect or correct the mistake, leading to a significantly higher error rate, particularly in categories related to schema understanding and table manipulation. It’s desirable for future work to explore agentic methods with multiple rounds of API calls.